0.4p-3.2=4.8p+4.2-0.7p
solve for p

Answer by funmath(2933)

0.4p-3.2=4.8p+4.2-0.7p
0.4p-3.2=4.1p+4.2
0.4p-0.4p-3.2=4.1p-0.4p+4.2
-3.2=3.7p+4.2
-3.2-4.2=3.7p+4.2-4.2
-7.4=3.7p
-7.4/3.7=3.7p/3.7
-2=p
:
Hint-If you really hate decimals, you can eliminate them by multiplying everything by 10.
10(0.4p-3.2)=10(4.8p+4.2-0.7p)
4p-32=48p+42-7p
4p-32=41p+42
4p-4p-32=41p-4p+42
-32=37p+42
-32-42=37p+42-42
-74=37p
-74/37=37p/37
-2=p
Happy Calculating!!!